A probability distribution has a mean of 35 and a standard deviation of 3. Use Chebychevs inequality to find a bound on the probability that an outcome of the experiment lies between the following (a) 30 and 40 at least 0.64 % (b) 25 and 45 at least 0.91 X% Need Help? Read Watch
Answer #1

Proposition Let X be a random variable having finite mean μ and finite variance σ2. Let k e R (i.e., k is a strictly positive real number). Then, the following inequality, called Chebyshevs inequality, holds:

(a)

Now X is less than 30 or greater than 40 if and only if,

| X - mu| geq k

where, mu = 35 and 5. So the probability that X <= 30 or X >= 40 is,

32 425 25

So X will be between 30 and 40 with probability = 1 - (9/25) = 16/25 = 64% (not 0.64% as written in the answer box)

(b)

Now X is less than 25 or greater than 45 if and only if,

| X - mu| geq k

where, mu = 35 and k 10. So the probability that X <= 25 or X >= 45 is,

32 2 102 100

So X will be between 30 and 40 with probability = 1 - (9/100) = 91/100 = 91% (not 0.91% as written in the answer box)

** If this answer does not match please comment